OpenStudy (anonymous):

What is the solution set for |3x – 24| less than or equal to 27?

OpenStudy (anonymous):

\[\left[ -3,21 \right]\]

OpenStudy (anonymous):

these are my choices x greater than or equal to 1 or x less than or equal to 17 x less than or equal to –1 and x greater than or equal to 17 x less than or equal to 17 or x less than or equal to –17 x greater than or equal to –1 and x less than or equal to 17

OpenStudy (precal):

\[\left| 3x-24 \right|\le27\] \[-27\le3x-24\le27\] add 24 to both sides \[-3\le3x\le51\] divide by 3 to both sides \[-1\le x \le17\] this means all numbers from -1 to 17

OpenStudy (precal):

or x is greater than or equal to -1 and 17 is greater than or equal to x